# Latent Diffusion for Neural Spiking Data

### v1

@@ 1 Introduction
For high-dimensional neural recordings, analyses that infer low-dimensional structures have been very useful for making sense of such data [11]. For example, latent variable models (LVMs) are often used to identify neural population dynamics not apparent at the level of single neurons [61, 32, 40]. More recently, deep learning-based approaches based on variational autoencoders (VAEs) [25, 44] have become particularly popular for inferring latent neural representations due to their expressiveness and ability to scale to large, heterogeneous neural recordings with behavioral covariates  [51, 36, 63, 16, 46].
However, in addition to learning latent representations, another important consideration is the ability to act as faithful generative models of the data. In other words, models should be able to produce diverse, realistic samples of the neural activity they were trained on, ideally in a behavior- or stimulus-dependent manner. Models with such capabilities not only afford better interpretability analyses and diagnoses for whether structures underlying the data are accurately learned, but have a variety of downstream applications surrounding the design of closed-loop in silico experiments. For example, with faithful generative models, one can simulate population responses to hypothetical sensory, electrical, or optogenetic stimuli, as well as possible neural activity underlying hypothetical movement patterns.
Most VAE-based approaches focus on the interpretability of the inferred latents, but not the ability to generate realistic and diverse samples when conditioning on external covariates, while sample-realistic models (e.g., based on generative adversarial networks (GANs) [17]) do not provide access to underlying low-dimensional representations. As such, there is a need for models of neural population spiking activity that both provide low-dimensional latent representations and can (conditionally) generate realistic neural activity.
@@ 2.2 Regularized autoencoder for neural spiking data
To prevent the autoencoder from predicting highly localized Poisson rates, which have sharp peaks at input spike locations, we further regularize training using coordinated dropout [23], i.e., we randomly mask input spikes and compute the loss on the predicted rates at the masked locations (see appendix A1.2 for details).
@@
So far, LDNS (like most state-space models for neural data) uses a Poisson observation model, which assumes that all statistical dependencies are mediated by the latent state. To address this limitation and to capture dynamics and variability, which are “private” to individual neurons (such as refractory periods or burstiness), we propose to learn an autoregressive observation model. We make the predicted Poisson rates for each neuron $i$ dependent also on recent spiking history, by including additional spike history couplings $h_{i}$ [41, 32], resulting in the observation model
@@
where ${T^{\prime}}$ corresponds to the time-lagged window length. This modification is learned post hoc, and the parameters $h_{i}$ are fitted with a maximum-likelihood approach. See appendix A1.3 for details.
@@ 2.3 Denoising Diffusion Probabilistic Models
Diffusion models may be conditioned on fixed-length and time-varying covariates $\mathbf{y}$, in which case we learn the approximate reverse transition $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$. For details on the conditioning mechanisms, see appendix A1.4.
@@ 3.1 Datasets and tasks
We first evaluate the performance of LDNS on a synthetic spiking dataset where we have access to the ground-truth firing rates and latents. We choose the Lorenz attractor [28] as a low-dimensional, non-linear dynamical system commonly used in neuroscience [7, 36]. We simulate rates as an affine mapping from the 3-dimensional system to a 128-dimensional neural space, and sample from a Poisson distribution to generate spiking data. Next, we showcase the applicability of LDNS on two neural datasets: We apply our method on a highly complex dataset of human neural activity (128 units) recorded during attempted speech [56]. This dataset poses a challenge to many modeling approaches due to the different imagined sentences, resulting in variable lengths of the neural time series (between 2-10 seconds with a sampling rate of 50 Hz). Finally, we apply LDNS to model premotor cortical activity (182 units) recorded from monkeys performing a delayed center-out reach task with barriers [9, 38]. The multi-modal nature of the dataset allows us to assess both unconditional as well as conditional generation of neural spiking activity given monkey reach directions and entire velocity profiles of the performed reaches. For training details, see appendix A2.
@@
Baselines    We compare LDNS to the most commonly known VAE-based latent variable model: Latent Factor Analysis via Dynamical Systems (LFADS [51, 36, 47]), which has been shown to outperform various classical latent variable models on a variety of tasks ([38], see appendix A3 for details). To ensure that we use optimal hyperparameters for LFADS, we follow the auto-ML pipeline proposed by Keshtkaran et al. [24]. This approach, termed AutoLFADS, has been shown to perform better than the original LFADS on benchmark tasks [38].
@@
Metrics   For all experiments, we assess how well LDNS-generated samples match the spiking data in the training distribution. Concretely, we compare population-level statistics by computing 1) the distribution over the population spike count, which sums up all spikes co-occurring in the population in a single time bin (i.e., spike count histogram), and 2) pairwise correlations of LDNS samples and the spiking data for each pair of neurons. For single-neuron statistics, we compare 3) the mean and 4) standard deviation of the inter-spike-interval distribution for each neuron (mean isi and std isi). When multiple spikes occur in a single time bin, the spike times are distributed equally in this bin [12]. All metrics are calculated equally for LDNS, AutoLFADS, and the real data.
@@ 3.2 LDNS captures the true spiking data distribution with an underlying Lorenz system
Figure 2: Realistic generation of spiking data with underlying chaotic dynamics. a) Synthetic spiking data from an underlying Lorenz system. b) Accurate, smooth rate predictions of the autoencoder for held-out spiking data. c) Plotted trace of sampled latents (256 bins training length, left) and 16×16$\times$16 × the original training length (middle). The sampled latent distribution matches the PSD of the autoencoder latents (right; median, 10%, and 90% percentiles). d) LDNS population spike count histogram, and pairwise cross-correlation match the training distribution. e) LDNS single neuron statistics, i.e., mean inter-spike interval (isi) and std isi, match the training distribution.
We simulate trials of length 256 timesteps from the three-dimensional (chaotic) Lorenz system (Fig. 2a). The regularized autoencoder extracts smooth latent time series (eight latent dimensions) from the 128-dimensional spiking data, resulting in smooth firing rate predictions that closely match the ground-truth rates (Fig. 2b, Supp. Fig. A5 A6). We then train a diffusion model on the extracted autoencoder latents. Latents sampled from the diffusion model (red) preserve the attractor geometry of the Lorenz system (Fig. 2c, left, three of the eight latent dimensions), indicating that LDNS preserves a meaningful latent space. The architectural choice of S4 layers allows for length generalization: although we train on time segments of 256-time steps, we can sample and successfully generate latent trajectories that are much longer, but still accurately reflect the Lorenz dynamics (Fig. 2c, middle, 16×16$\times$16 × longer generation). Overall, the latent time series distribution is captured well by the diffusion model, with matching power spectral densities (PSD) per latent dimension (Fig. 2e, other dimensions in Supp. Fig. A6). To assess the sampling fidelity of the generated synthetic neural activity, we compute a variety of spike statistics frequently used in neuroscience. LDNS captures both population-level statistics, such as the population spike count histogram and pairwise correlations between neurons (Fig. 2d), as well as single-neuron statistics, quantified by the mean and standard deviation of inter-spike-intervals (Fig. 2e). These results demonstrate that LDNS can both perform inference of low-dimensional latents and provide high-fidelity diffusion-based generation that perfectly captures the statistics of the ground-truth synthetic data.
@@ 3.3 Modeling variable-length trials of neural activity recorded in human cortex
Figure 3: Unconditional generation of variable-length trials of human spiking data during attempted speech. a) Multi-unit activity is recorded from speech production-related regions of the brain (top) during attempted vocalization of variable-length sentences (bottom). b) Neural activity during sentences of different lengths. c) LDNS unconditionally sampled trials with different lengths. d) LDNS population spike count histogram, and mean and std of the isi match those of the data. e) Correlation matrices of the data (left) and LDNS samples (middle), and scatterplot of the pairwise correlations of data vs. LDNS samples (right).
@@
This approach allows us to infer time-aligned latents underlying the cortical activity of the participants, compressing the population activity by a factor of four before training an unconditional diffusion model on these latents. Resulting samples of LDNS, mimicking the human cortical activity, are visually indistinguishable from the real data (Fig. 3b,c). This is also reflected in closely matched population spike count histograms (Fig. 3d, left), and single neuron statistics such as mean and standard deviation of the inter-spike interval (Fig. 3d, right). While LDNS tends to overestimate some pairwise correlations, it captures prominent features of the correlation structure in the data (Fig. 3e, Pearson correlation coefficient $r=0.47$).
@@ 3.4 Realistic generation of spiking data from a monkey performing reach tasks
Figure 4: Realistic generation of spiking data in a monkey performing reach tasks. a) A monkey performs diverse reach movements in different mazes. b) Neural activity during a reach trial and a sampled trial from LDNS. c) The LDNS population spike count histogram, and pairwise correlations match those of the data. d) LDNS mean- and std isi match the monkey data distribution. e) Auto-correlation of data, LDNS samples with Poisson observations (left), and LDNS samples with spike history, grouped according to correlation strength.
We further evaluate LDNS in a different setting by applying it to model sparse spiking data recorded from a monkey performing a reaching task constrained by barriers that form a maze (Fig. 4a, left). The variety of different maze architectures leads to diverse reach movements of both curved and straight reaches (Fig. 4a, right).
We again infer low-dimensional latent trajectories that capture the shared variability of the neural population and then train an unconditional diffusion model on these latents. Sampled spikes from LDNS closely resemble the true, sparse population data (Fig. 4b), and closely match population-level spike statistics (Fig. 4c). Single neuron statistics in this low spike count regime (a maximum of three spikes per neuron in 5 ms bins) are also captured well and on par with or better than our baseline comparison model, AutoLFADS [36, 24]. For a quantitative comparison of population and single neuron statistics, see Tab. 1 and Supp. Fig. A3.
Table 1: Model metrics comparison. $D_{KL}$ for the population spike count histogram and RMSE comparisons. Mean and standard deviation across 5 folds sampled with replacement.
@@
To overcome this mismatch, we train an additional spike history-dependent autoregressive observation model on top of the inferred rates (LDNSsh, for spike history). In contrast to the Poisson samples, autoregressive samples can capture this aspect of neural spiking data very accurately while also improving the overall fit to the empirical auto-correlation (Fig. 4e, right). Moreover, the post hoc optimization of these filters also improves modeling of other single-neuron, as well as population-level statistics, such as the population spike count histogram or the mean of the isi (Tab. 1, Supp. Fig. A3).
Poisson LDNS allows for the generation of spiking data that is on par or better in terms of sampling fidelity than previous approaches. Incorporating spike-history dependence and sampling spikes autoregressively allows us to further increase the realism of generated spike trains, leading to a large improvement on several of the considered metrics.
@@ 3.5 Conditional generation of neural activity given reach directions or velocity profiles
Finally, an even more challenging task that is intriguing for hypothesis generation is the ability to mimic an entire experiment and ask what the neural activity would have looked like if the monkey had performed a particular hypothetical movement. To this end, we train a diffusion model on the same autoencoder-inferred latents but now condition on entire velocity traces (Fig. 5d). Velocity-conditioned LDNS is able to produce different samples of neural activity that are consistent with, but not exact copies of, the reach trajectories of the held-out trials given as the conditioning covariate. Such closed-loop conditioning experiments open the possibility of making predictions about neural activity during desired unseen behaviors, and thus make experimentally testable predictions.
In summary, LDNS not only produces faithful spiking samples on this dataset but also allows for flexible conditioning.
Figure 5: Generation conditioned on monkey reach directions and velocity traces. a) Closed loop assessment: do conditionally generated latents translate to neural activity consistent with the desired direction or reach movement? b) Unseen reach movements (data) and corresponding movements decoded from the rates predicted by the autoencoder (ae). c) Decoded reach directions of LDNS samples conditioned on initial reach angles $\theta$. d) Decoded reach directions of LDNS samples conditioned on 3 unseen reach movements (velocities $v_{x},v_{y}$)

### v2

@@ 1 Introduction
For high-dimensional neural recordings, analyses that infer low-dimensional structures have been very useful for making sense of such data [11]. For example, latent variable models (LVMs) are often used to identify neural population dynamics not apparent at the level of single neurons [61, 32, 40]. More recently, deep learning-based approaches based on variational autoencoders (VAEs) [26, 44, 51, 36, 63, 22, 6, 16] have become particularly popular for inferring latent neural representations due to their expressiveness and ability to scale to large, heterogeneous neural recordings with behavioral covariates  [36, 63, 16].
However, in addition to learning latent representations, another important consideration is the ability to act as faithful generative models of the data. In other words, models should be able to produce diverse, realistic samples of the neural activity they were trained on, ideally in a behavior- or stimulus-dependent manner. Models with such capabilities not only afford better interpretability analyses and diagnoses for whether structures underlying the data are accurately learned, but have a variety of downstream applications surrounding the design of closed-loop in silico experiments. For example, with faithful generative models, one can simulate population responses to hypothetical sensory, electrical, or optogenetic stimuli, as well as possible neural activity underlying hypothetical movement patterns. Most VAE-based approaches focus on the interpretability of the inferred latents, but not the ability to generate realistic and diverse samples when conditioning on external covariates, while sample-realistic models (e.g., based on generative adversarial networks (GANs) [17]) do not provide access to underlying low-dimensional representations. As such, there is a need for models of neural population spiking activity that both provide low-dimensional latent representations and can (conditionally) generate realistic neural activity.
@@ 2.2 Regularized autoencoder for neural spiking data
To prevent the autoencoder from predicting highly localized Poisson rates, which have sharp peaks at input spike locations, we further regularize training using coordinated dropout [24], i.e., we randomly mask input spikes and compute the loss on the predicted rates at the masked locations (details in appendix A1.2).
@@
So far, LDNS (like most latent variable models for neural data) uses a Poisson observation model, which assumes that all statistical dependencies are mediated by the latent state. To address this limitation and to capture dynamics and variability, which are “private” to individual neurons (such as refractory periods or burstiness), we propose to learn an autoregressive observation model. We make the predicted Poisson rates for each neuron $i$ dependent also on recent spiking history, by including additional spike history couplings $h_{i}$ [41, 32], resulting in the observation model
@@
where ${T^{\prime}}$ corresponds to the time-lagged window length. This modification is learned post hoc, and the parameters $h_{i}$ are fit with a maximum-likelihood objective (details in appendix A1.3). This approach does not alter the latent dynamics, while augmenting the model with single-neuron autoregressive dynamics. We observe that including spike history increases the realism of generated data and enables us to accurately capture single-neuron autocorrelation structures (Sec. 3.4).
@@ 2.3 Denoising Diffusion Probabilistic Models
Diffusion models may be conditioned on fixed-length and time-varying covariates $\mathbf{y}$, in which case we learn the approximate reverse transition $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$. Details on the conditioning mechanisms in appendix A1.4.
@@ 3.1 Datasets and tasks
We first evaluate the performance of LDNS on a synthetic spiking dataset where we have access to the ground-truth firing rates and latents. We choose the Lorenz attractor [29] as a low-dimensional, non-linear dynamical system commonly used in neuroscience [7, 36]. We simulate rates as an affine mapping from the 3-dimensional system to a 128-dimensional neural space, and sample from a Poisson distribution to generate spiking data. Next, we showcase the applicability of LDNS on two neural datasets: We apply our method on a highly complex dataset of human neural activity (128 units) recorded during attempted speech [56]. This dataset poses a challenge to many modeling approaches due to the different imagined sentences, resulting in variable lengths of the neural time series (between 2-10 seconds with a sampling rate of 50 Hz). Finally, we apply LDNS to model premotor cortical activity (182 units) recorded from monkeys performing a delayed center-out reach task with barriers [9, 38]. The multi-modal nature of the dataset allows us to assess both unconditional as well as conditional generation of neural spiking activity given monkey reach directions and entire velocity profiles of the performed reaches. See appendix A2,A3 for data and training details.
For the unconditional generation of monkey reach recordings (Sec. 3.4), we train both a Poisson observation model as well as a spike history-dependent autoregressive observation model. For all other experiments, we only train a Poisson observation model.
Figure 2: Realistic generation of spiking data with underlying chaotic dynamics. a) Synthetic spiking data from an underlying Lorenz system with a Poisson observation model. b) Accurate, smooth rate predictions of the autoencoder for held-out spiking data. c) Plotted trace of sampled latents (256 bins training length, left) and 16×16$\times$16 × the original training length (middle). The sampled latent distribution matches the PSD of the autoencoder latents (right; median, 10%, and 90% percentiles). d) LDNS population spike count histogram (kde: kernel density estimate) and pairwise cross-correlations match the training distribution. e) LDNS single neuron statistics, i.e., mean inter-spike interval (isi) and std isi, match the training distribution.
@@
Baselines    We compare LDNS to the most commonly known VAE-based latent variable model: Latent Factor Analysis via Dynamical Systems (LFADS [51, 36, 47]), which has been shown to outperform various classical latent variable models on a variety of tasks ([38], details in appendix A4). To ensure that we use optimal hyperparameters for LFADS, we follow the auto-ML pipeline proposed by Keshtkaran et al. [25]. This approach, termed AutoLFADS, has been shown to perform better than the original LFADS on benchmark tasks [38]. For the unconditional generation of monkey reach recordings, we further compared to additional VAE baselines [21, 62] (appendix A5).
@@
Metrics   For all experiments, we assess how well LDNS-generated samples match the spiking data in the training distribution. Concretely, we compare population-level statistics by computing 1) the distribution over the population spike count, which sums up all spikes co-occurring in the population in a single time bin (i.e., spike count histogram), and 2) pairwise correlations of LDNS samples and the spiking data for each pair of neurons. For single-neuron statistics, we compare 3) the mean and 4) standard deviation of the inter-spike-interval distribution for each neuron (mean isi and std isi). When multiple spikes occur in a single time bin, the spike times are distributed equally in this bin [12]. To further evaluate population dynamics, we compare the principal components of smoothed spikes.
@@ 3.2 LDNS captures the true spiking data distribution with an underlying Lorenz system
We simulate trials of length 256 timesteps from the three-dimensional (chaotic) Lorenz system (Fig. 2a). The regularized autoencoder extracts smooth latent time series (eight latent dimensions) from the 128-dimensional spiking data, resulting in smooth firing rate predictions that closely match the ground-truth rates (Fig. 2b, Supp. Fig. A2,A3). We then train a diffusion model on the extracted autoencoder latents. Latents sampled from the diffusion model (red) preserve the attractor geometry of the Lorenz system (Fig. 2c, left, three of the eight latent dimensions), indicating that LDNS preserves a meaningful latent space. The architectural choice of S4 layers allows for length generalization: although we train on time segments of 256-time steps, we can sample and successfully generate latent trajectories that are much longer, but still accurately reflect the Lorenz dynamics (Fig. 2c, middle, 16×16$\times$16 × longer generation). In comparison, LFADS exhibits instabilities when generating such longer sequences (appendix  A6.1). Overall, the latent time series distribution is captured well by the diffusion model, with matching power spectral densities (PSD) per latent dimension (Fig. 2c, right, other dimensions in Supp. Fig. A3).
To assess the sampling fidelity of the generated synthetic neural activity, we compute a variety of spike statistics frequently used in neuroscience. LDNS captures both population-level statistics, such as the population spike count histogram and pairwise correlations between neurons (Fig. 2d), as well as single-neuron statistics, quantified by the mean and standard deviation of inter-spike-intervals (Fig. 2e). LDNS also captures the temporal correlation structure of the data (Supp. Fig. A4). These results demonstrate that LDNS can both perform inference of low-dimensional latents and provide high-fidelity diffusion-based generation that perfectly captures the statistics of the ground-truth synthetic data.
@@ 3.3 Modeling variable-length trials of neural activity recorded in human cortex
Figure 3: Unconditional generation of variable-length trials of human spiking data during attempted speech. a) Multi-unit activity is recorded from speech production-related regions of the brain (top) during attempted vocalization of variable-length sentences (bottom). b) Neural activity during sentences of different lengths. c) LDNS unconditionally sampled trials with different lengths, using the Poisson observation model. d) LDNS population spike count histogram, and mean and std of the isi match those of the data. e) Correlation matrices of the data (left) and LDNS samples (middle), and scatterplot of the pairwise correlations of data vs. LDNS samples (right).
@@
This approach allows us to infer time-aligned latents underlying the cortical activity of the participants, compressing the population activity by a factor of four before training an unconditional diffusion model on these latents. Resulting samples of LDNS, mimicking human cortical activity, are visually indistinguishable from the real data (Fig. 3b,c, additional samples in Supp. Fig. A7). This is reflected in closely matched population spike count histograms (Fig. 3d, left), and single neuron statistics such as mean and standard deviation of the inter-spike interval (Fig. 3d, right). Additionally, real and LDNS-sampled spikes display similar population dynamics, as reflected in the top principal components (Supp. Fig. A8). While LDNS tends to overestimate some pairwise correlations, it captures prominent features of the correlation structure in the data (Fig. 3e, Pearson correlation coefficient $r=0.47$), and our analysis indicates that this slight mismatch already arises at the autoencoder stage (Supp. Fig. A9).
@@ 3.4 Realistic generation of spiking data from a monkey performing reach tasks
Figure 4: Realistic generation of spiking data in a monkey performing reach tasks. a) A monkey performs diverse reach movements in different mazes. b) Neural activity during a reach trial and a sampled trial from LDNS with a Poisson observation model. c) The LDNS population spike count histogram, and pairwise correlations match those of the data. d) LDNS mean- and std isi match the monkey data distribution. e) Auto-correlation of data, LDNS samples with Poisson observations (left), and LDNS samples with spike history, grouped according to correlation strength.
We further evaluate LDNS in a different setting by applying it to model sparse spiking data recorded from a monkey performing a reaching task constrained by barriers that form a maze (Fig. 4a, left). The variety of different maze architectures leads to diverse reach movements of both curved and straight reaches (Fig. 4a, right). We again infer low-dimensional latent trajectories that capture the shared variability of the neural population and then train an unconditional diffusion model on these latents. Sampled spikes from LDNS closely resemble the true, sparse population data (Fig. 4b, additional samples in Supp. Fig. A11), and closely match population-level spike statistics (Fig. 4c). Single neuron statistics in this low spike count regime (a maximum of three spikes per neuron in 5 ms bins) are also captured well (Fig. 4d), and are on par with or better than LFADS [36, 25] (see Table 1 for summary of main comparisons, and appendix A5 for additional baselines [21, 62]). Beyond spiking statistics, we observe that LDNS also preserves the temporal structure of population dynamics, as reflected in the top principal components of smoothed spikes (Supp. Fig. A15). Thus, LDNS can generate spiking data that is faithful at the level of both single-neuron and population dynamics.
Table 1: Model metrics comparison. $D_{KL}$ for the population spike count histogram and RMSE comparisons. Mean and standard deviation across 5 folds sampled with replacement. sh represents observation models with spike history. Bolded entries represent best-performing values for Poisson and spike-history observation models.
@@
LDNS $\mathbf{0.0039\pm 3.9\text{e-}4}$ $\mathbf{0.0025\pm 1.1\text{e-}5}$ $\mathbf{0.037\pm 0.001}$ $\mathbf{0.023\pm 0.001}$
@@
To overcome this mismatch, we train an additional spike history-dependent autoregressive observation model on top of the inferred rates (LDNSsh, for spike history). In contrast to the Poisson samples, autoregressive samples can capture this aspect of neural spiking data very accurately while also improving the overall fit to the empirical auto-correlation (Fig. 4e, right). Moreover, the post hoc optimization of these filters also improves modeling of other single-neuron, as well as population-level statistics, such as the population spike count histogram or the mean of the isi (Table 1, Supp. Fig. A13).
We view this post-hoc augmentation as a key modular contribution, which can be flexibly applied to other generative models. To this end, we extend AutoLFADS with spike history dependence (LFADSsh), improving its performance across metrics. The augmented LFADSsh also captures the dip in autocorrelation at 5 ms lags (Supp. Fig. A12). Still, in both observation model variants, LDNS maintains superior or comparable performance (Table 1).
Thus, Poisson LDNS allows for the generation of spiking data that is on par or better in terms of sampling fidelity than previous approaches. Incorporating spike-history dependence and sampling spikes autoregressively allows us to further increase the realism of generated spike trains, leading to a large improvement on several of the considered metrics.
Figure 5: Generation conditioned on monkey reach directions and velocity traces. a) Closed loop assessment: do conditionally generated latents translate to neural activity consistent with the desired direction or reach movement? b) Unseen reach movements (data) and corresponding movements decoded from the rates predicted by the autoencoder (ae). c) Decoded reach directions of LDNS samples conditioned on initial reach angles $\theta$. d) Decoded reach directions of LDNS samples conditioned on 3 unseen reach movements (velocities $v_{x},v_{y}$). e) Straight reaches from the test set used for velocity conditioning. f) LDNS sampled latents conditioned on trajectories shown in e) vary smoothly over time and reflect information about reach angles. g) PCs of sampled LDNS latents shown in f) reveal meaningful and separable information about behavior.
@@ 3.5 Conditional generation of neural activity given reach directions or velocity profiles
An even more challenging task that is intriguing for hypothesis generation is the ability to mimic an entire experiment and ask what the neural activity would have looked like if the monkey had performed a particular hypothetical movement. To this end, we train a diffusion model on the same autoencoder-inferred latents but now condition on entire velocity traces (Fig. 5d). Velocity-conditioned LDNS is able to produce different samples of neural activity that are consistent with, but not exact copies of, the reach trajectories of the held-out trials given as the conditioning covariate. Such closed-loop conditioning experiments open the possibility of making predictions about neural activity during desired unseen behaviors, and thus make experimentally testable predictions.
Finally, to understand how LDNS incorporates behavioral information, we analyzed latent trajectories that were conditionally sampled based on straight reach movements in different directions (Fig. 5e). Individual samples of latent trajectories vary smoothly within a trial (Fig. 5f), while reach direction varies smoothly across samples in the first principal component (PC1) of the latents (Fig. 5g, left). Projection onto the first two PCs of latent trajectories shows clear clustering by reach direction (Fig. 5g, right), and we show that such clustering arises already at the autoencoder stage (Supp. Fig. A17).
In summary, LDNS not only produces faithful spiking samples but also allows for flexible conditioning. Furthermore, LDNS learns an interpretable latent space with behaviorally-relevant structure.